# Scene Informer: Anchor-based Occlusion Inference and Trajectory Prediction in Partially Observable Environments

### 16 Feb 2024 v2

@@ III Scene Informer
Reasoning over observed and occluded agents that might interfere with a planned trajectory is critical to ensure the safety of the AV maneuvers. It involves processing a range of vectorized input modalities provided by upstream perception frameworks, such as observed agents and lane graphs while taking into account the partial observability of the environment. We propose an environment prediction framework called Scene Informer (see Fig. 2) that takes in observed trajectories of other agents, vectorized maps, and infers the true state of the occlusion interest and the future motion of observed agents. Our framework consists of transformer encoder and decoder pairing. The encoder takes in vectorized representation of the observed agents and maps to create the scene embeddings [40, 41]. The decoder receives the scene embeddings and reasons in terms of spatial anchors for both observed agents and occlusions. Spatial anchors are created for each observed agent and randomly populated in the occlusion of interest. It allows intelligent queries for occlusions that might interact with the planned trajectory rather than inferring all occlusions at once, which is computationally expensive and unnecessary. Finally, for each provided anchor, the decoder infers a probability of occupancy and the most likely trajectories with their corresponding probabilities to model the multi-future behavior of the occluded agents.
@@ III-A Problem Statement and Input Representation
Given an environment state representation that consists of observed agent information and road map information , our objective is to infer the occlusion and predict future trajectories from the perspective of the ego vehicle. Agent information represents agents over history horizon with features consisting of position, heading, velocity, and agent’s dimensions. Road map information is denoted as which describes polylines with points with feature vector . Each point is represented by its position, vector to the next point in a polyline, and polyline type (e.g. crosswalk, lane boundary, stop sign, etc.) We represent the occlusion to infer and agents to predict by a set of spatial anchors , where is the number of anchors defined by their position. The anchors are uniformly spread in the occluded area and assigned to the current position of the observed agents as shown in Fig. 3. For each anchor, we predict most likely trajectories over the prediction horizon with probability distribution , and probability of the anchor being occupied .
@@ III-B Framework
We use a transformer architecture that has been shown to excel at modeling multimodal and unstructured inputs of varying sizes [40, 41, 13]. Our proposed framework consists of scene encoder and anchor decoder. The scene encoder processes unstructured multimodal input, such as past agent trajectories and vectorized maps, and generates scene embeddings. The anchors that define agent and occlusions to predict are provided to the anchor decoder which cross-attends them with the scene embeddings to generate anchor embeddings. Those embeddings are then used to predict the probability of the anchor being occupied, and top future trajectories with their corresponding probability distribution.
Scene encoder: It aggregates the state representation that includes agent information and road map information modalities. Each modality type is modeled with a dedicated PointNet polyline encoder layer [42, 13] as it allows us to encode polylines of varying lengths:
where is an input for each modality with points of dimensionality , MLP is a multi-layer perceptron network, is a max-pooling operator over the last feature dimension, and is an encoded polyline. For different agent types, we have a dedicated set of parameters for MLP. Subsequently, all polylines are concatenated to form . They are provided to the transformer encoder which allows us to aggregate input modalities with a varying number of tokens:
@@
Training loss: Our loss is a mix of regression, trajectory classification, and anchor occupancy classification components. For a given anchor , if it is occupied, we identify the predicted mode out of with a hard-assignments strategy [10, 43]. Subsequently, we maximize the log-likelihood associated with the selection of mode j, the generation of the ground truth trajectory by mode , and the occupation of anchor . Conversely, if anchor remains unoccupied, the log-likelihood is maximized for its unoccupied state. The cumulative loss is computed by aggregating these components over all anchors:
@@ IV-B Implementation details
Scene encoder: We implement four polyline encoders for cars, bikes, pedestrians, and maps. Polyline encoders output a set of feature vectors with a size of 256 for each polyline. We use a transformer encoder with four layers, four heads, 256 hidden size, and a 2048 feedforward dimension.
@@ IV-C Experimental Setup and Metrics
Our evaluation metrics encompass the classification of occupancy for unobserved agents and a regression analysis comparing predicted trajectories to the ground truth for both observed and unobserved agents. In the classification task, we assess the effectiveness in inferring which anchors are occupied within occlusions, focusing on the classification accuracy for occupied and free anchors (/). For the regression task, we evaluate the alignment between our predicted trajectories and the ground truth using minimum average and final displacement errors (/).
We evaluate the performance on occluded agents by benchmarking against existing occlusion inference methods, such as K-means PaS [7], GMM PaS [8], and MAVOI [8]. None of the open-source approaches is capable of inferring future trajectories of occluded agents. To evaluate the inferred trajectory, we compare our method with variations of our framework: the vanilla trajectory prediction, akin to recent works of [41, 13, ], and an occlusion inference adaptation that focuses exclusively on reasoning over occlusions. Even though the vanilla trajectory prediction assumes full observability, we can still query it to infer the trajectory of occluded agents by providing an anchor with the ground truth position of the occluded agent. For the observed agents, we compare Scene Informer with a vanilla trajectory prediction trained on fully observable data and limited observability data. The latter lets us determine whether we can enhance the robustness of vanilla trajectory prediction to partial observability without explicitly incorporating occlusion inference reasoning. We train all occlusion-based models on a fully observable dataset with agents in a single occlusion removed for inference purposes. We then assess the robustness of all models by evaluating them under various observability scenarios. This involves using a dataset where the likelihood of an occlusion being generated by an agent varies. Contrary to the training dataset, it applies to all agents potentially creating multiple occlusions in the scene. It ranges from (full observability), with increments of , , and (partial observability), to (limited observability) (see Fig. 3).
@@ V Discussion
Occlusion inference performance: We report the numerical evaluation of the occlusion inference performance of our approach compared to the baselines and the impact of evaluation observability assumptions in Table I. Our method consistently outperforms all baselines in both occupied and free accuracy metrics, regardless of the visibility setting, often by a substantial margin. The improvement is in the range of - and - points on occupied and free cells, respectively. Notably, as the visibility assumption in the evaluation dataset gets more restrictive, our model tends to predict more occupied cells—a potentially desirable trait for safety-critical applications. On the regression task, Scene Informer outperforms a vanilla trajectory prediction baseline and a variation trained only on the occluded agents. As the observation becomes more limited in evaluation, the performance gap between models, particularly when compared to vanilla trajectory prediction, widens in favour of our approach. In the limited observability setting, the of Scene Informer is less than that of the vanilla trajectory prediction, which underscores the ability of our framework to robustly reason about occlusions across a range of observability scenarios. In addition, our framework outperforms its variation trained solely on occluded agents, which implies that occlusion inference and trajectory prediction are complementary tasks and should be trained together.
@@
Figure 4 provides examples of inferred occlusions. As shown, our approach effectively predicts occupancy and the future trajectory based on other agents’ motions and road layouts. The predicted occupancy map realistically reflects the true motion of partially observable traffic, infers potential unobserved agents and assigns a higher probability of occupancy to actively used parts of drivable spaces. Figure 5 shows how the observed motion of the other traffic participants impacts the occlusion inference. We modify the observations of agents to indicate potentially different flow of traffic. We demonstrate that Scene Informer adapts its predictions realistically in response to changes in the behavior of observed agents.

### 09 Mar 2024 v3

@@ III Scene Informer
Reasoning over observed and occluded agents that might interfere with a planned trajectory is critical to ensure the safety of the AV maneuvers. It involves processing a range of vectorized input modalities provided by upstream perception frameworks, such as observed agents and lane graphs while taking into account the partial observability of the environment. We propose an environment prediction framework called Scene Informer (see Fig. 2) that takes in observed trajectories of other agents, vectorized maps, and infers the true state of the occlusion interest and the future motion of observed agents. Our framework consists of transformer encoder and decoder pairing. The encoder takes in vectorized representation of the observed agents and maps to create the scene embeddings [40, 41]. The decoder receives the scene embeddings and reasons in terms of spatial anchors for both observed agents and occlusions. For observed agents, spatial anchors indicate the most recent position. In the case of occlusions, they represent points randomly placed within the occlusion of interest. It allows intelligent queries for occlusions that might interact with the planned trajectory rather than inferring all occlusions at once, which is computationally expensive and unnecessary. Finally, for each provided anchor, the decoder infers a probability of occupancy and the most likely trajectories with their corresponding probabilities to model the multi-future behavior of the observed and potentially present but occluded agents.
@@ III-A Problem Statement and Input Representation
Given an environment state representation that consists of observed agent information and road map information , our objective is to infer the occlusion and predict future trajectories from the perspective of the ego vehicle. Agent information represents agents over history horizon with features consisting of position, heading, velocity, and agent’s dimensions. Road map information is denoted as which describes polylines with points with feature vector . Each point is represented by its position, vector to the next point in a polyline, and polyline type (e.g. crosswalk, lane boundary, stop sign, etc.) We represent the occlusion to infer and agents to predict by a set of spatial anchors , where is the number of anchors defined by their position. anchors are assigned to the most recent positions of the observed agents, while anchors are uniformly distributed within the occluded areas, as shown in Fig. 2. For each anchor, we predict most likely trajectories over the prediction horizon with probability distribution , and probability of the anchor being occupied .
@@ III-B Framework
We use a transformer architecture that has been shown to excel at modeling multimodal and unstructured inputs of varying sizes [40, 41, 13]. Our proposed framework consists of scene encoder and anchor decoder. The scene encoder processes multimodal input, such as past agent trajectories and vectorized maps, and generates scene embeddings. The anchors that define agent and occlusions to predict are provided to the anchor decoder which cross-attends them with the scene embeddings to generate anchor embeddings. Those embeddings are then used to predict the probability of the anchor being occupied, and top future trajectories with their corresponding probability distribution.
Scene encoder: It aggregates the state representation that includes agent information and road map information modalities. Each agent type is encoded with a dedicated multi-layer perceptron, and each road polyline is modeled with a PointNet polyline encoder layer [42, 13] as it allows us to encode polylines of varying lengths:
where MLP is a multi-layer perceptron network, and is a max-pooling operator over the last feature dimension. For different agent types, we have a dedicated set of parameters for MLP. Subsequently, all representations are concatenated to form . They are provided to the transformer encoder which aggregates input modalities with a varying number of tokens:
@@
Training loss: Our loss is a mix of regression, trajectory classification, and anchor occupancy classification components. For a given anchor , if it is occupied, we identify the predicted mode out of possible trajectories with a hard-assignments strategy [10, 43]. Subsequently, we maximize the log-likelihood associated with the selection of mode j, the generation of the ground truth trajectory by mode , and the occupation of anchor . Conversely, if anchor remains unoccupied, the log-likelihood is maximized for its unoccupied state. The cumulative loss is computed by aggregating these components over all anchors:
@@ IV-B Implementation details
Scene encoder: We implement three MLPs for cars, bikes, pedestrians, and a polyline encoder for maps. Each output a set of feature vectors with a size of 256 for each observed agent and each polyline. We use a transformer encoder with four layers, four heads, 256 hidden size, and a 2048 feedforward dimension.
@@ IV-C Experimental Setup and Metrics
We assess our framework’s performance in terms of occlusion inference of unobserved agents and trajectory prediction of observed agents in a partially observable environment. Our evaluation metrics encompass the classification of occupancy for unobserved agents and a regression analysis comparing predicted trajectories to the ground truth for observed and unobserved agents. In the classification task, we assess the effectiveness in inferring which anchors are occupied within occlusions, focusing on the classification accuracy for occupied and free anchors (/). For the regression task, we evaluate the alignment between our predicted trajectories and the ground truth using minimum average and final displacement errors (/).
We evaluate the performance on occluded agents by benchmarking against existing occlusion inference methods, such as K-means PaS [7], GMM PaS [8], and MAVOI [8]. None of the open-source approaches is capable of inferring future trajectories of occluded agents. To evaluate the inferred trajectory, we compare our method with variations of our framework: the vanilla trajectory prediction that reasons only about observed agents, and an occlusion inference adaptation that focuses exclusively on reasoning over occlusions. Even though the vanilla trajectory prediction assumes full observability, we can still query it to infer the trajectory of occluded agents by providing an anchor with the ground truth position of the occluded agent. For the observed agents, we compare Scene Informer with a vanilla trajectory prediction variations of our framework trained on fully observable data and limited observability data, respectively. The latter lets us determine whether we can enhance the robustness of vanilla trajectory prediction to partial observability without explicitly incorporating occlusion inference reasoning. We then assess the robustness of all models by evaluating them under various observability scenarios. This involves using a dataset where the likelihood of an occlusion being generated by an agent varies. Contrary to the training dataset, it applies to all agents potentially creating multiple occlusions in the scene. It ranges from (full obs.), with increments of , , and (partial obs.), to (limited obs.) (see Fig. 3).
@@ V Discussion
Occlusion inference performance: We report the evaluation of the occlusion inference performance of our approach compared to the baselines and the impact of evaluation observability assumptions in Table I. Our method consistently outperforms all baselines in both occupied and free accuracy metrics, regardless of the visibility setting. The improvement is in the range of - and - points on occupied and free cells, respectively. Notably, as the visibility assumption in the evaluation dataset gets more restrictive, our model tends to predict more occupied cells—a potentially desirable trait for safety-critical applications.
On the regression task, Scene Informer outperforms a vanilla trajectory prediction baseline and a variation trained only on the occluded agents. As the observation becomes more limited, the performance gap between models, particularly when compared to vanilla trajectory prediction, widens in favour of our approach. In the limited observability setting, the of Scene Informer is less than that of the vanilla trajectory prediction, which underscores the ability of our framework to robustly reason about occlusions across a range of observability scenarios. In addition, our framework outperforms its variation trained solely on occluded agents, which implies that occlusion inference and trajectory prediction are complementary tasks.
@@
Figure 4 provides examples of inferred occlusions. As shown, our approach effectively predicts occupancy and the future trajectory based on other agents’ motions and road layouts. The predicted occupancy map realistically reflects the true motion of partially observable traffic, infers potential unobserved agents and assigns a higher probability of occupancy to actively used parts of drivable spaces. Figure 5 shows how the observed motion of the traffic participants impacts the occlusion inference. We modify the observations to indicate potentially different flow of traffic. We demonstrate that Scene Informer adapts its predictions realistically in response to changes in the behavior of observed agents.